OpenStudy (devonhoward15):

Jermiah is factoring 0.64t^24-9 by using a^2-b^2=(a+b)(a-b). Which expression will he use for the value of a?

OpenStudy (devonhoward15):

@Bobo-i-bo

OpenStudy (devonhoward15):

i think 0.64

OpenStudy (bobo-i-bo):

Why?

OpenStudy (devonhoward15):

yahoo lol

OpenStudy (bobo-i-bo):

Well yahoo is very wrong :P

OpenStudy (devonhoward15):

thought so lol

OpenStudy (bobo-i-bo):

So \(0.64t^{24}-9\) is in the form of the difference of two squares, i.e. it is the square of one number... let's say \(a\), minus the square of another number which we'll call \(b\). In other words it is in the form \(a^2-b^2\). So you have to find what \(a\) is. Following me? Do you understand better/see what you have to do?

OpenStudy (devonhoward15):

0.64t^24-x^2

OpenStudy (bobo-i-bo):

?

OpenStudy (devonhoward15):

idk lol

OpenStudy (devonhoward15):

i was trying to find a

OpenStudy (bobo-i-bo):

If I tell you that 0.64t^2-9 is in the form of a^2-b^2, then a^2=0.64t^24 and b^2=9

OpenStudy (bobo-i-bo):

Basically the square of a is equal to 0.64t^24 and the square of b is equal to 9, so what is b and what is a?

OpenStudy (devonhoward15):

a=0.64t^24 and b=9

OpenStudy (bobo-i-bo):

nono, the SQUARE of a is equal to 0.64t^24 and the SQUARE of b is equal to 9. a is not equal to 0.64t^24 and b is not equal to 9

OpenStudy (bobo-i-bo):

So basically \(a^2=0.64t^{24}\) and \(b^2=9\)

OpenStudy (devonhoward15):

ohh ok

OpenStudy (devonhoward15):

i just forgot to square a and b

OpenStudy (bobo-i-bo):

so what's a and what is b?

OpenStudy (devonhoward15):

am i supposed to put 0.64t^24 in my calculator?

OpenStudy (devonhoward15):